OpenStudy (anonymous):

What is the distance from the line x+y=-7 to the point (4, -1) ?

OpenStudy (anonymous):

I think it should be done with the dot product but it seems that I forgot how to do it :(, lets google!

OpenStudy (anonymous):

do you think google can answer this :((((((

OpenStudy (anonymous):

google knows everything :-) Found a formula already: rearranging the line : x+y+7=0 (4,-1) D=modulus(4*1+(-1)*1+7)/sqrt(1^2 +1^2)= =10/sqrt2

OpenStudy (anonymous):

Slope of line is -1 so slope of perpendicular is 1 Get a line with that slope through (4,-1)

OpenStudy (anonymous):

@kris100 is that u mr aringo?

OpenStudy (anonymous):

@nudnick Why ?

OpenStudy (anonymous):

@nudnick who are you?